The lenght of a rectangle is 3 inches more than its width.If the area of the rectangle is 238 square inches,what is its lenght?
Answer by checkley71(8403)

L=W+3
THUS LW=238
(W+3)W=238
W^2+3W-238=0
(W+17)(W-14)=0
W-14=0
W=1
THEN L=14+3
L=17
PROOF
17*14=238
238=238